OpenStudy (anonymous):

Determine whether the ratios are proportional. 1. 16/18, 18/24 2. 2/3, 10/15

OpenStudy (shamil98):

reduce the fractions. if they are the same when reduced they are proportional .

OpenStudy (shamil98):

16/18 = 4/9 18/24= 3/4 1 is not proportional

OpenStudy (shamil98):

8/9* thanks zimmah